I need explanation and solution:
The product of four consecutive natural numbers is 255024. Find these numbers.

Answer by rothauserc(4718)

let x be the first natural number, then
x * (x+1) * (x+2) * (x+3) = 255024
(x^2+x)* (x^2+5x+6) = 255024
x^4 + 6x^3 + 11x^2 + 6x - 255024 = 0
we get two sets of solutions for this problem
(21, 22, 23, 24) or (-24, -23, -22, -21)
check
21 * 22 * 23 * 24 = 255024
answer checks